OpenStudy (anonymous):

f(x) = -15ln(33x) Compute f '(x), then find the exact value of f ' (4).

OpenStudy (anonymous):

If f(x) -= ln (u), then f'(x) = du/u The derivative of ln(33x) = 1/(33x) times 33 dx = dx/x So your answer is -15/x Hence, f'(4) = -15/4